OpenStudy (bananas1234):

Factor: x2 – 7x + 12

OpenStudy (hockeychick23):

ok so you need to find 2 numbers that multiply to equal 12 and add to equal 7. Do you know what two numbers will give you that result?

OpenStudy (bananas1234):

-4 x -3 = 12 and -4 + -3 + -7

OpenStudy (hockeychick23):

correct so you get (x-4)(x-3)

OpenStudy (bananas1234):

X = 4 and X = 3?

OpenStudy (hockeychick23):

when you factor you don't solve for x but yes 3 and 4 are the two numbers you will use because x*x= x^2 x*-3= -3x -4*x= -4x -4*-3= 12 added together you get x^2-3x-4x+12 which simplified is x^2-7x+12 your factored equation will look like this: (x-4)(x-3)

OpenStudy (bananas1234):

i see thank you!!!

OpenStudy (hockeychick23):

sure :)